OpenStudy (anonymous):

A 60 kg student in a rowboat on a still lake decides to dive off the back of the boat. The student's horizontal acceleration is 2.0 m/s2 while in contact with the boat. What horizontal force does the student exert on the boat?

OpenStudy (anonymous):

16 a = 18e3/ 0.15 = 120,000 m/s2 17 m = 60 /10 = 6 kg g’ at moon = 10 /6 = 1.6 m/s² 18 120 N toward the front of the boat 19 30 times the ball's weight => 30times g = 290 m/s² 20 After leaving the dog's hair…

OpenStudy (anonymous):

@Azonto do you need a little help? @S_CG Force = mass * acceleration = 60 * 2 = 120N